A=P(1+r)t P=1250, r=4%, t=10years
Answer by checkley77(12844)

A=P(1+r)^t P=1250, r=4%, t=10years
A=1,250(1+.04)^10
A=1,250(1.04)^10
A=1,250*1.48
A=1,850 ANS.